# Chat with AI: The Surprising Turn of Real-time Video Communication from Human to AI

### v1

@@ Abstract.
AI Video Chat emerges as a new paradigm for Real-time Communication (RTC), where one peer is not a human, but a Multimodal Large Language Model (MLLM). This makes interaction between humans and AI more intuitive, as if chatting face-to-face with a real person. However, this poses significant challenges to latency, because the MLLM inference takes up most of the response time, leaving very little time for video streaming. Due to network uncertainty and instability, transmission latency becomes a critical bottleneck preventing AI from being like a real person. To address this, we propose Artic, an AI-oriented Real-time Communication framework, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. To reduce bitrate dramatically while maintaining MLLM accuracy, we propose Context-Aware Video Streaming that recognizes the importance of each video region for chat and allocates bitrate almost exclusively to chat-important regions. To avoid packet retransmission, we propose Loss-Resilient Adaptive Frame Rate that leverages previous frames to substitute for lost/delayed frames while avoiding bitrate waste. To evaluate the impact of video streaming quality on MLLM accuracy, we build the first benchmark, named Degraded Video Understanding Benchmark (DeViBench). Finally, we discuss some open questions and ongoing solutions for AI Video Chat.
@@ 1. Introduction
AI Video Chat is a new paradigm for Real-time Communication (RTC). Since OpenAI released GPT-4o (Hurst et al., 2024), Multimodal Large Language Models (MLLMs) have been continuously emerging, such as Qwen2.5-Omni (Xu et al., 2025), Gemini-1.5 (Team et al., 2024), VITA-1.5 (Fu et al., 2025), and OmniLive (Zhang et al., 2024). Compared to traditional LLMs, MLLMs enable users to directly input video and audio for interaction, rather than just text. This makes the interaction between humans and AI more intuitive as if chatting face-to-face with a real person (Chen et al., 2024b). However, MLLMs require high-performance computing devices to support real-time inference (such as 8*A100 GPUs (Qiu et al., [n. d.])). Mobile devices (like phones or smart glasses) cannot meet the computing requirements, which makes MLLMs inevitably deployed in the cloud. So in existing systems, the client sends user video and audio to the cloud for MLLM inference, and the cloud then feeds back responses to users, as shown in Figure 1.
AI Video Chat raises significant challenges to RTC transmission latency. To ensure a fluent interactive experience, the end-to-end response latency of video chat needs to remain below 300 ms (Lai et al., 2022). In traditional video chat, the human peer on the other side can respond instantly. Therefore, transmission latency accounts for the vast majority of end-to-end response latency. To reduce transmission latency, current state-of-the-art RTC frameworks (such as WebRTC (Carlucci et al., 2016)) adopt technologies like Adaptive Bitrate (ABR) (Mao et al., 2017; Yan et al., 2020; Akhtar et al., 2018; Huang et al., 2014; Jiang et al., 2012), Congestion Control (Cardwell et al., 2017; Dong et al., 2015, 2018; Meng and Xu, 2024), and Forward Error Correction (FEC) (Meng et al., 2024; An et al., 2025; Rudow et al., 2023; Cheng et al., 2024; Li et al., 2023) to satisfy user experience. However, in AI Video Chat, responses are generated through MLLM in an autoregressive manner, which is time-consuming. Even when inputting only audio tokens, the computational latency is at least 232 ms (Hurst et al., 2024). To constrain the response latency below 300 ms, the time left for transmission is at most 68 ms, which is difficult to guarantee. So the large latency makes users clearly feel that the other side is not a real person.
@@
To realize this vision, we propose Artic, an AI-oriented Real-time Communication framework, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. We begin by recognizing the main differences between AI Video Chat and traditional RTC: First, QoE changes from human perceptual quality to MLLM response accuracy. Second, the receiver throughput is far lower than the sender throughput. Third, uplink is more pressing than Downlink. Then, through prototype measurements, we identify two key factors for low latency: ultra-low bitrate and loss resilience. Based on three insights, we make contributions:
@@
  • Frame Rate is FEC (§3.3). Since MLLMs process videos at a very low frame rate, most received frames are redundant. Although redundant frames cause bitrate waste, they can also substitute for lost/delayed frames. So we propose Loss-Resilient Adaptive Frame Rate to simultaneously minimize bitrate waste and packet retransmission.
@@ 2.1. Main differences between AI Video Chat and traditional RTC
QoE changes from human perceptual quality to MLLM response accuracy. In traditional RTC, QoE focuses on measuring the perceptual quality of human eyes. For example, using penalty terms like stalling time (Cheng et al., 2024; Yan et al., 2020) or quality variance (Chen et al., 2024a) to measure temporal stability. Using SSIM (Yan et al., 2020) or VMAF (vma, 2025) to measure visual quality. However, in AI Video Chat, the viewer of the video changes from humans to MLLMs. At this point, most perception-related metrics are no longer needed, and the optimization objective of QoE becomes the accuracy and latency of MLLM’s responses. So RTC strategy can undergo a significant turn. For example, to reduce latency, temporal stability (such as frequent bitrate adjustments) and visual quality (like lowering bitrate in Figure 3) can be sacrificed, as long as accuracy is enough.
The receiver throughput is far lower than the sender throughput. In traditional RTC, the data throughput at the receiver is comparable to that at the sender, for example, both are 1920*1080 resolution at a 30 FPS frame rate. However, in AI Video Chat, MLLM is limited by context length (finite number of tokens) and cannot fully process the received video. Therefore, the received video needs to be actively downsampled before being fed to the MLLM. In terms of frame rate, existing AI Video Chat systems support a maximum processing rate of only 2 FPS (Xu et al., 2025; Wang et al., 2025b; Fu et al., 2025). In terms of resolution, regardless of how high the original resolution is, it will be downsampled to no more than 602,112 pixels (Xu et al., 2025).
@@ 2.2. What factors affect the transmission latency of AI Video Chat?
To analyze the factors affecting transmission latency in AI Video Chat, we build a prototype and conduct preliminary measurements. Specifically, we implement a WebRTC-based unidirectional video transmission system and a network emulator. Under given bandwidth (10 Mbps) and one-way network delay (30 ms), we run video transmission for a total duration of 40,489 seconds, and collect statistics on frame latency (the time difference from the frame being sent to being completely received) with different packet loss rates and bitrates, as shown in Figure 2. The results indicate:
First, when the bitrate exceeds the bandwidth, frame latency becomes enormous. This is because excessive bitrate causes congestion, where packet accumulation causes latency to increase rapidly. Therefore, existing RTC systems employ ABR algorithms to set the bitrate as close as possible to (but below) the bandwidth, maximizing video quality while avoiding stalling, as shown in the grey region of Figure 2. Second, when the bitrate does not exceed the bandwidth, frame latency also increases as the bitrate increases. This is due to the fact that each packet has a limited size (around 1400 bytes). A higher bitrate means each frame will be divided into more packets. Due to packet loss, more packets mean the probability of each frame being completely received in one attempt decreases. For packets that are not received, retransmission may be required, leading to increased latency. Therefore, even when the bitrate is below the bandwidth, AI Video Chat can further reduce the bitrate to achieve lower latency. This differs from traditional ABR and offers another space for bitrate selection, as shown in the yellow region of Figure 2. Third, when the packet loss rate increases, frame latency increases. Similarly, this is also caused by more retransmissions.
@@ 2.3. Key Insights and Potential Gains
Video should be Context-Aware. According to §2.2 reducing video bitrate can decrease transmission latency. To reduce bitrate, existing methods typically increase quantization parameters or lower resolution (Schwarz et al., 2007), which inevitably degrades video quality. Interestingly, the degradation in video quality does not necessarily lead to a decrease in MLLM accuracy, which depends on the current chat context. As illustrated in Figure 3, when the user asks ”Could you tell me the present score of the game?”, even if the video bitrate is reduced from 4000 Kbps to 200 Kbps, the MLLM can still answer accurately. However, when the user asks ”What logo is seen on the jersey of the player covering his mouth?”, the blurry video leads to incorrect responses. This is because, in different chat contexts, the MLLM needs to focus on different video regions. Meanwhile, different video regions are affected differently by low bitrate. Thus, rather than reducing bitrate in a context-agnostic manner, the video should be context-aware. More bitrate should be allocated to chat-important regions, while less bitrate should be allocated to chat-irrelevant regions.
Figure 3. Why video should be context-aware in AI Video Chat. In the first dialogue, even if the video bitrate decreases from 4000 Kbps to 200 Kbps, the MLLM can still response accurately. But in the second dialogue, the blurry video leads to incorrect responses. Therefore, rather than reducing bitrate in a context-agnostic manner, bitrate allocation should be determined by the current chat context (§2.3).
@@
Figure 4. Frame Rate is FEC. MLLM processes video at a very low frame rate (green region), so higher frame rates will introduce redundancy (red region). Higher frame rates on one hand cause bitrate waste, while on the other hand enhance loss resilience (§2.3).
Frame Rate is FEC. According to §2.2, besides bitrate, packet loss rate also has an important impact on latency. We find that frame rate can make trade-offs between bitrate and packet loss rate in existing AI Video Chat systems. As described in §2.1, MLLM processes videos at very low frame rates (such as 2 FPS), while the client still transmits videos at conventional frame rates (like 30 FPS). After receiving the video, MLLM will sample frames at a uniform time interval (such as 500 ms) for processing, while the remaining frames are not used. This causes most received video frames to be redundant. So to reduce video bitrate, the most reasonable solution is to use the same low frame rate (2 FPS) at the sender. Figure 4 shows the relationship between frame rate and video bitrate. It can be seen that lowering the frame rate can reduce bitrate while not affecting MLLM accuracy.
Figure 6. How to achieve loss resilience? When the expected frame is not completely received on time, the MLLM directly discards it and uses the previous redundant frame, without waiting for retransmission.
In contrast to reducing bitrate, increasing the frame rate can enhance resilience to packet loss. As illustrated in Figure 6, rather than sending video at 2 FPS, the client can send at a higher frame rate. When the expected frame (e.g., sampled at 500 ms intervals) is not received on time, the MLLM can directly discard it and use a previous (redundant) frame, without waiting for retransmission. In this way, redundant frames can handle packet loss in an FEC manner. This presents a trade-off: higher frame rates increase redundancy for better loss tolerance, but also lead to higher bitrates. So the frame rate should be selected based on loss rate (§3.3).
@@
3. Towards RTC for AI
For the above insights, we begin by constructing the first benchmark evaluating how video quality affects MLLM accuracy, named Degraded Video Understanding Benchmark (DeViBench). Then we design two modules: context-aware video streaming and loss-resilient adaptive frame Rate.
@@ 3.1. DeViBench
Video Preprocessing. To allow MLLMs to understand the quality degradation caused by low bitrates, we transcode the original videos to low-bitrate versions (200 Kbps). The low bitrate video and the original video are horizontally concatenated into one video. Then this concatenated video will be input to the MLLM for understanding and QA generation.
QA Generation. To enable MLLMs to generate QA pairs based on the concatenated video, we carefully designed a prompt with guidance from persona, context, core task, execution steps, constraints, and output format, as shown in Figure 8. This prompt ensures that MLLMs can recognize quality differences and generate quality-sensitive QA pairs.
QA Filtering. The generated QA pairs will be filtered. We separately input the original video and the low bitrate video into the MLLM and use the generated QA pairs for questioning. If the answer from the original video is correct and the answer from the low bitrate video is wrong, we accept this QA pair. In practice, 25.2% of the QA pairs can be accepted.
Cross Verification. Since the answer generated by MLLM may also be incorrect, this cannot be filtered out through the above testing. Hence, we utilize another MLLM for cross-verification. We feed the above accepted question into another MLLM, and if the new answer is consistent with the above accepted answer, we finally approve this QA pair. In our experiments, 57.7% of the accepted QA pairs can pass cross-verification. Considering all the above validations together, finally 14.6% of the generated QA pairs are valid.
@@ 3.2. Context-Aware Video Streaming
In this section, we describe how to achieve context-aware streaming, significantly reducing bitrate while maintaining MLLM accuracy. According to §2.3, we first leverage the CLIP model to compute the semantic correlation between user words and video regions. To ensure real-time computing on mobile devices, we adopt Mobile-Clip (Vasu et al., 2024). Specifically, given the current user words $\mathcal{T}$ and the latest video frame $F\in\mathbb{R}^{H\times W\times 3}$, we first partition $F$ into non-overlapping patches $\left\{P_{mn}\mid 1\leq m\leq\left\lfloor H/N\right\rfloor,1\leq n\leq\left\lfloor W/N\right\rfloor\right\}$, where each patch $P_{mn}\in\mathbb{R}^{N\times N\times 3}$ represents a video region. Then, the CLIP visual encoder $\phi_{v}(\cdot):\mathbb{R}^{N\times N\times 3}\rightarrow\mathbb{R}^{d}$ is employed to extract patch-wise features $f_{mn}^{v}=\phi_{v}\left(P_{mn}\right)$, while the CLIP language encoder $\phi_{l}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$ encodes the user words $\mathcal{T}$ into a semantic features $f^{l}=\phi_{l}(\mathcal{T})$. Here $d$ denotes the unified feature dimension. Semantic correlation $\rho_{mn}$ between user words and patches are then computed as cosine similarities:
@@
Where $\gamma$ is the temperature parameter, which we set 3 to aggressively penalizes irrelevant regions ($\rho_{mn}\ll 1$).
We evaluate the performance gains in Figure 9. The results demonstrate that context-aware streaming can dramatically lower the bitrate while maintaining MLLM accuracy. For example, when the bitrate is reduced from 800 Kbps to 400 Kbps (50% reduction), the MLLM (Qwen2.5-Omni (Xu et al., 2025)) accuracy drops from 0.73 to 0.33. After integrating context-aware streaming, the accuracy only decreases from 0.93 to 0.87.
3.3. Loss-Resilient Adaptive Frame Rate
In this section, we describe how to select the frame rate based on the current loss rate, thereby jointly reducing redundancy and packet retransmission. According to Figure 6, when the expected frame is not received, Artic uses a previous redundant frame as the substitute. So given the loss rate, we aim to minimize the frame rate while maximizing the probability that at least one redundant frame is received on time.
Specifically, let $p_{t}\in[0,1]$ denote the packet loss rate during decision epoch $t$ (simplified to i.i.d. when the decision epoch is sufficiently small). Let $\kappa_{t}\in\mathbb{Z}^{+}$represent the average number of packets per frame within epoch $t$. The probability of successfully receiving a frame is then given by:
(3) $P_{f}=\left(1-p_{t}\right)^{\kappa_{t}}.$
Let $R_{M}$ denote the inference frame rate of the MLLM and let $R_{t}$ represent the video frame rate. So the MLLM samples one frame from every $K=$ $\left\lceil R_{t}/R_{M}\right\rceil$ frames. The probability of successfully receiving at least one frame in $K$ frames is:
(4) $P_{s}=1-(1-P_{f})^{K}$
To avoid retransmissions, it should satisfy $P_{s}\geq(1-\epsilon)$, where $\epsilon=0.001$ is the residual term. Therefore, we finally select the minimum frame rate $R_{t}$:
(5) $R_{t}\geq R_{M}\cdot\left\lceil\frac{\ln\epsilon}{\ln\left(1-\left(1-p_{t}\right)^{\kappa_{t}}\right)}\right\rceil$
We evaluate the performance gains in Figure 10. The preliminary results show that our method can reduce the latency caused by stalling from 25.3 ms to 1.2 ms, while avoiding the bitrate waste brought by high frame rates.
Figure 10. Eliminating the latency caused by stalling, while avoiding the waste brought by high frame rates.

### v2

@@ Abstract.
AI Video Chat emerges as a new paradigm for Real-time Communication (RTC), where one peer is not a human, but a Multimodal Large Language Model (MLLM). This makes interaction between humans and AI more intuitive, as if chatting face-to-face with a real person. However, this poses significant challenges to latency, because the MLLM inference takes up most of the response time, leaving very little time for video streaming. Due to network uncertainty, transmission latency becomes a critical bottleneck preventing AI from being like a real person. To address this, we call for AI-oriented RTC research, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. We begin by recognizing the main differences between AI Video Chat and traditional RTC. Then, through prototype measurements, we identify that ultra-low bitrate is a key factor for low latency. To reduce bitrate dramatically while maintaining MLLM accuracy, we propose Context-Aware Video Streaming that recognizes the importance of each video region for chat and allocates bitrate almost exclusively to chat-important regions. To evaluate the impact of video streaming quality on MLLM accuracy, we build the first benchmark, named Degraded Video Understanding Benchmark (DeViBench). Finally, we discuss some open questions and ongoing solutions for AI Video Chat. DeViBench is open-sourced at: https://github.com/pku-netvideo/DeViBench.
Real-time Communication, Generative AI, Multimodal Large Language Model, AI Video Chat, Latency, Benchmark
journalyear: 2025copyright: acmlicensedconference: The 24th ACM Workshop on Hot Topics in Networks; November 17–18, 2025; College Park, MD, USAbooktitle: The 24th ACM Workshop on Hot Topics in Networks (HotNets ’25), November 17–18, 2025, College Park, MD, USAdoi: 10.1145/3772356.3772390isbn: 979-8-4007-2280-6/2025/11ccs: Computing methodologies Artificial intelligenceccs: Networks Application layer protocolsccs: Information systems Multimedia streamingccs: Human-centered computing Human computer interaction (HCI)
@@ 1. Introduction
AI Video Chat is a new paradigm for Real-time Communication (RTC). Since OpenAI released GPT-4o (Hurst et al., 2024), Multimodal Large Language Models (MLLMs) have been continuously emerging, such as Qwen2.5-Omni (Xu et al., 2025), Gemini-1.5 (Team et al., 2024), VITA-1.5 (Fu et al., 2025), and OmniLive (Zhang et al., 2024). Compared to traditional LLMs, MLLMs enable users to directly input video and audio for interaction, rather than just text. This makes the interaction between humans and AI more intuitive as if chatting face-to-face with a real person (Chen et al., 2024b). However, MLLMs require high-performance computing devices to support real-time inference (such as 8*A100 GPUs (Qiu et al., 2025)). Mobile devices (like phones or smart glasses) cannot meet the computing requirements, which makes MLLMs inevitably deployed in the cloud. So in existing systems, the client sends user video and audio to the cloud for MLLM inference, and the cloud then feeds back responses to users, as shown in Figure 1.
AI Video Chat raises significant challenges to RTC transmission latency. To ensure a fluent interactive experience, the response latency of video chat needs to remain below 300 ms (Lai et al., 2022). In traditional video chat, the human peer on the other side can respond instantly. Therefore, response latency largely stems from RTC’s end-to-end latency (including capture, transmission, decoding, and playback buffer latencies). Among them, transmission latency is sensitive to network conditions and continuously increases when the network deteriorates. To reduce transmission latency, current state-of-the-art RTC frameworks (such as WebRTC (Carlucci et al., 2016)) adopt technologies like Adaptive Bitrate (ABR) (Mao et al., 2017; Yan et al., 2020; Akhtar et al., 2018; Huang et al., 2014; Jiang et al., 2012), Congestion Control (Cardwell et al., 2017; Dong et al., 2015, 2018; Meng and Xu, 2024), and Forward Error Correction (FEC) (Meng et al., 2024; An et al., 2025; Rudow et al., 2023; Cheng et al., 2024; Li et al., 2023) to satisfy user experience. However, in AI Video Chat, responses are generated through MLLM in an autoregressive manner, which is time-consuming. Even when inputting only audio tokens, the computational latency is at least 232 ms (Hurst et al., 2024). To constrain the response latency below 300 ms, even ignoring other latencies in the RTC pipeline, the time left for transmission is at most 68 ms, which is difficult to guarantee. So the large latency makes users clearly feel that the other side is not a real person.
@@
To realize this vision, we call for AI-oriented RTC research, exploring the network requirement shift from ”humans watching video” to ”AI understanding video”. We begin by recognizing the main differences between AI Video Chat and traditional RTC: First, QoE changes from human perceptual quality to MLLM response accuracy. Second, jitter has no impact. Third, the receiver throughput is far lower than the sender throughput. Fourth, uplink is more pressing than Downlink. Then, through prototype measurements, we identify a key factor for low latency: ultra-low bitrate. Based on two insights, we make contributions:
@@ 2.1. Main differences between AI Video Chat and traditional RTC
QoE changes from human perceptual quality to MLLM response accuracy. In traditional RTC, QoE focuses on measuring the perceptual quality of human eyes. For example, using penalty terms like stalling time (Cheng et al., 2024; Yan et al., 2020) or quality variance (Chen et al., 2024a) to measure temporal stability. Using SSIM (Yan et al., 2020) or VMAF (vma, 2025) to measure visual quality. But in AI Video Chat, the viewer of the video changes from humans to MLLMs. At this point, most perception-related metrics are no longer needed, and the optimization objective of QoE becomes the accuracy and latency of MLLM’s responses. So the RTC strategy can undergo a significant turn. For example, to reduce latency, temporal stability (such as frequent bitrate adjustments) and visual quality (like lowering bitrate in Figure 4) can be sacrificed, as long as accuracy is enough.
Jitter has no impact. Due to network uncertainties (congestion, packet loss, etc.), even when the sender transmits frames at fixed time intervals, the time intervals of received frames will experience jitter. Direct playback will result in uneven video speed, causing stuttering. Therefore, traditional RTC employs a jitter buffer (Zhao et al., 2024), trading latency for smoothness. For MLLMs, the difference is that their perception of time does not rely on real physical time, but rather on positional encoding computation (Xu et al., 2025; Fu et al., 2025; Zhang et al., 2024). Positional encoding is only associated with the frame’s capture timestamp and unrelated to the actual receiving time; thus, jitter has no impact on the MLLM’s perception of the video. It means that in AI video chat, the buffer can be removed to reduce the latency.
The receiver throughput is far lower than the sender throughput. In traditional RTC, the data throughput at the receiver is comparable to that at the sender, for example, both are 1920*1080 resolution at a 30 FPS frame rate. However, in AI Video Chat, MLLM is limited by context length (finite number of tokens) and real-time inference, and cannot fully process the received video. Therefore, the received video needs to be actively downsampled before being fed to the MLLM. In terms of frame rate, existing AI Video Chat systems support a maximum processing rate of only 2 FPS (Xu et al., 2025; Wang et al., 2025b; Fu et al., 2025). In terms of resolution, regardless of how high the original resolution is, it will be downsampled to no more than 602,112 pixels (Xu et al., 2025). So traditional RTC contains massive redundancy that MLLMs cannot perceive, as shown in Figure 2.
Figure 2. MLLM processes video at a very low frame rate (green), so most frames are redundancy (red).
@@ 2.2. What factors affect the transmission latency of AI Video Chat?
To analyze the factors affecting transmission latency in AI Video Chat, we built a prototype and conducted preliminary measurements. Specifically, we implement a WebRTC-based unidirectional video transmission system and a network emulator. Under given bandwidth (10 Mbps) and one-way network delay (30 ms), we run video transmission for a total duration of 40,489 seconds, and collect statistics on transmission latency (the time from the frame being sent to being completely received, excluding the jitter buffer §2.1) with different packet loss rates and bitrates, as shown in Figure 3:
First, when the bitrate exceeds the bandwidth, transmission latency becomes enormous. This is because excessive bitrate causes congestion, where packet accumulation causes latency to increase rapidly. Therefore, existing RTC systems employ ABR algorithms to set the bitrate as close as possible to (but below) the bandwidth, maximizing video quality while avoiding stalling, as shown in the grey region of Figure 3. Second, when the bitrate does not exceed the bandwidth, transmission latency also increases as the bitrate increases. This is due to the fact that each packet has a limited size (around 1400 bytes). A higher bitrate means each frame will be divided into more packets. Due to packet loss, more packets mean the probability of each frame being completely received in one attempt decreases. For packets that are not received, retransmission may be required, potentially leading to increased latency. Therefore, even when the bitrate is below the bandwidth, AI Video Chat can further reduce the bitrate to achieve lower latency. This differs from traditional ABR and offers another space for bitrate selection, as shown in the yellow region of Figure 3.
@@ 2.3. Key Insights and Potential Gains
Video should be Context-Aware. According to §2.2 reducing video bitrate can decrease transmission latency. To reduce bitrate, existing methods typically increase quantization parameters (Schwarz et al., 2007), which inevitably degrades video quality. Interestingly, the degradation in video quality does not necessarily lead to a decrease in MLLM accuracy, which depends on the current chat context. As illustrated in Figure 4, when the user asks ”Could you tell me the present score of the game?”, even if the video bitrate is reduced from 4000 Kbps to 200 Kbps, the MLLM can still answer accurately. However, when the user asks ”What logo is seen on the jersey of the player covering his mouth?”, the blurry video leads to incorrect responses. This is because, in different chat contexts, the MLLM needs to focus on different video regions. Meanwhile, different video regions are affected differently by low bitrate. Thus, rather than reducing bitrate in a context-agnostic manner, the video should be context-aware. More bitrate should be allocated to chat-important regions, while less bitrate should be allocated to chat-irrelevant regions.
Figure 4. Why video should be context-aware in AI Video Chat. In the first dialogue, even if the video bitrate decreases from 4000 Kbps to 200 Kbps, the MLLM can still response accurately. But in the second dialogue from StreamingBench (Lin et al., 2024), the blurry video leads to incorrect responses. Thus, rather than reducing bitrate in a context-agnostic manner, bitrate allocation should be determined by the current chat context (§2.3).
@@
3. Towards RTC for AI: Case Study
We begin by constructing the first benchmark evaluating how video quality affects MLLM accuracy, named Degraded Video Understanding Benchmark (DeViBench). Then we present a case study: Context-Aware Video Streaming.
@@ 3.1. DeViBench
Video Preprocessing. To allow MLLMs to understand the quality degradation caused by low bitrates, we transcode the original videos to low-bitrate versions (200 Kbps) using x265 from ffmpeg version N-118035-gc1e3d55f99. The low bitrate video and the original video are horizontally concatenated into one video. Then this concatenated video will be input to the MLLM for understanding and QA generation.
QA Generation. To enable MLLMs to generate QA samples based on the concatenated video, we carefully designed a prompt with guidance from persona, context, core task, execution steps, constraints, and output format, as shown in Figure 7. This prompt ensures that MLLMs can recognize quality differences and generate quality-sensitive QA samples. To facilitate judging whether the answer is correct, we generate multiple-choice questions with four options (A, B, C, D). Users can directly calculate accuracy by matching the answer letters, without needing to measure semantic consistency. We also encourage MLLMs to generate questions that require at least multiple frames to answer, in order to enhance the temporal dependency of the questions. Qwen3-VL-plus thinking (qwe, 2025) is adopted as the generator.
@@
QA Filtering. The generated QA pairs will be filtered. We separately input the original video and the low bitrate video into the MLLM and use the generated QA pairs for questioning. If the answer from the original video is correct and the answer from the low bitrate video is wrong, we accept this QA pair. In practice, Qwen2.5-Omni (Xu et al., 2025) is adopted as the filter and 11.16% of the QA pairs can be accepted.
Cross Verification. Since the answer generated by MLLM may also be incorrect, this cannot be filtered out through the above testing. Hence, we utilize another MLLM for cross-verification. We feed the above accepted question into another MLLM, and if the new answer is consistent with the above accepted answer, we finally approve this QA pair. In our experiments, GLM-4.5V thinking (Team et al., 2025) is adopted as the verifier and 70.61% of the accepted QA pairs can pass cross-verification. Considering all the above validations together, finally 7.8% of the generated QA pairs are valid.
Finally, we produce 1,074 QA samples, with details summarized in Table 1, including QA types, total duration, total money spent, and total time cost. We also analyze the distribution of different QA types, as shown in Figure 8. In terms of categories, there are text-rich understanding (54.84%), action perception (17.03%), attribute perception (14.43%), counting (6%), object perception (5.9%), and spatial understanding (1.8%). In terms of temporal dependency, 34.45% of the questions necessitate multiple frames for answering, whereas 65.55% are answerable with a single frame. To confirm whether these MLLM-generated QA samples are usable, we spot-check 100 QA samples for manual answering. Among them, 95% of the generated questions are answerable by humans, and 84% of the generated answers are correct.
Table 1. Benchmark summary
Number of QA samples 1,074
QA sample types 6*2
Total duration (s) 180,000
Total money spent (\$) 68.47
Total time cost (s) 99,471
Figure 8. Distribution of our generated QA samples. Outer ring: QA categories. Inner ring: Whether the question requires multiple frames to answer.
@@ 3.2. Context-Aware Video Streaming
In this section, we describe how to achieve context-aware streaming, significantly reducing bitrate while maintaining MLLM accuracy. According to §2.3, we first leverage the CLIP model to compute the semantic correlation between user words and video regions. To ensure real-time computing on mobile devices, we adopt Mobile-Clip (Vasu et al., 2024). Specifically, given the current user words $\mathcal{T}$ and the latest video frame $F\in\mathbb{R}^{H\times W\times 3}$, we first partition $F$ into non-overlapping patches $\left\{P_{mn}\mid 1\leq m\leq\left\lfloor H/N\right\rfloor,1\leq n\leq\left\lfloor W/N\right\rfloor\right\}$, where each patch $P_{mn}\in\mathbb{R}^{N\times N\times 3}$ represents a video region. Then, the CLIP visual encoder $\phi_{v}(\cdot):\mathbb{R}^{N\times N\times 3}\rightarrow\mathbb{R}^{d}$ is employed to extract patch-wise features $f_{mn}^{v}=\phi_{v}\left(P_{mn}\right)$, while the CLIP language encoder $\phi_{l}(\cdot):\mathcal{T}\rightarrow\mathbb{R}^{d}$ encodes the user words $\mathcal{T}$ into a semantic features $f^{l}=\phi_{l}(\mathcal{T})$. Here $d$ denotes the unified feature dimension. Semantic correlation $\rho_{mn}$ between user words and patches is then computed as cosine similarities:
@@
Where $\gamma$ is the temperature coefficient, set to 3 to aggressively penalize irrelevant regions ($\rho_{mn}\ll 1$). To achieve fine-grained QP control, we adopt H.265 implemented by Kvazaar (Viitanen et al., 2016) to encode ours and baseline. Except for the QP values, ours and baseline use the same encoding parameters. The frame rate is consistent with the video source (e.g., 60 FPS). The specific Kvazaar command lines can be found in our open-source link. As for decoding, both ours and baseline adopt x265 from decord 0.6.0, maintaining the same decoding parameters. We test with Qwen2.5-Omni (Xu et al., 2025). Code and model are frozen before testing. We keep the default settings (the same random seed, system prompt, and configuration as officially recommended), without tuning for the QA samples.
Figure 10. An example of accuracy gain. We visualize the frame input to the MLLM. (a) Encoded with default settings. (b) Encoded with CLIP-informed QP. (c) The CLIP-informed QP map. The results show that even with similar bitrates (430 Kbps vs. 425 Kbps), ours allocates more bits to chat-important regions (e.g., purple circles) and fewer bits to chat-irrelevant regions (e.g., yellow circles), thus improving MLLM accuracy.
We evaluate the performance gains  [^1] Since DeViBench is continuously scaling up and iterating, the experiments in Figure 9 are frozen at an earlier version (small-scale, free-response, also available in the open-source link), rather than the version reported in §3.1. This is because video encoding takes a long time to run, and time constraints prevented us from rerunning the experiments on the current version. During Kvazaar encoding, the target bitrate often differs greatly from the actual bitrate. So we use a trial-and-error approach to ensure that the actual bitrates of ours and the baseline are comparable. Each video requires many encoding iterations, causing large time costs. Despite not updating the experiments, we speculate that the baseline’s accuracy would be higher on the current version. This is because, compared to the previous free-response questions, multiple-choice questions are easier to answer. On one hand, the options, as part of the question prompt, provide sufficient hints to the MLLMs. On the other hand, even if the video is too blurry to see clearly, MLLMs can still make a vague guess from the ABCD options (with at least 25% accuracy). in Figure 9. The results demonstrate that context-aware streaming can dramatically lower the bitrate while maintaining MLLM accuracy. For example, when the bitrate is reduced from 827.9 Kbps to 426.4 Kbps (48.5% reduction), the MLLM accuracy drops from 0.73 to 0.33. After integrating context-aware streaming, as the bitrate drops from 850.1 Kbps to 432.7 Kbps, the accuracy only decreases from 0.93 to 0.87. To intuitively demonstrate the benefits, we visualize two sampled frames fed into the MLLM, as shown in Figure 10. The results show that even with similar bitrates (430 Kbps vs. 425 Kbps), our method allocates more bits to chat-important regions (e.g., purple circles) and fewer bits to chat-irrelevant regions (e.g., yellow circles), thus improving MLLM accuracy.